# A is for Absorption: Studying Feature Splitting and Absorption in Sparse Autoencoders

### v4

@@ 5 Related work
5.1 Applications of Probes and SAEs for Model Interpretability
Probing methods have often been used to extract interpretable information from language models. However, the existence of such a representation does not guarantee that the model relies on this representation in its computation graph (Elazar et al., 2021).
Prior work has shown that many human-interpretable concepts in LLM activations are represented as linear directions in activation space, known as the linear representation hypothesis (Elhage et al., 2022; Park et al., 2024). Li et al. (2023) used non-linear probes to recover board representations from a transformer trained on Othello scripts (“OthelloGPT”). However, Nanda (2023) later showed that linear representations were not only recoverable but also editable.
Recent work has focused on applying SAEs to extract human-interpretable explanations of model internals. Karvonen et al. (2024) investigated how SAEs represent board states of Chess and Othello, and attempted to develop a more objective way of evaluating SAEs. However, it is difficult to apply this to existing SAEs trained on real LLM activations.
Other work has noted poor recall / precision of SAE latents compared to known proxies (Olah et al., 2024b; Kissane et al., 2024; Templeton et al., 2024). We build on this work by evaluating a large number of Gemma Scope SAEs to demonstrate how precision / recall is mediated by sparsity, and also offer a possible explanation of low recall due to feature absorption.
5.2 Character-level information in language models
The ability of LLMs to learn character-level information from ostensibly character-blind tokens has been studied by various scholars, though no clear mechanism has yet been established. Kaushal & Mahowald (2022) trained MLPs from the embedding layers of GPT-J as probes for each letter in the alphabet, again finding good performance that implied character-level information was represented, but did not look into the model internals to explain how these representations were being used. In a follow-up work, Watkins & Bloom (2023) demonstrated that even linear probes on the embedding layers perform comparably well to MLPs in extracting character-level information.
In contrast to the above approaches, we train various probes for each character on multiple layers, and compare with SAE latents for a variety of layers.
5.3 Decomposing SAE Latents
The phenomenon of SAEs of different sides splitting a feature into various smaller latents was first described in Bricken et al. (2023), which noted that different SAE widths and sparsities induce latents of different granularity, with wider SAEs often learning more specific variants of features.
Bussmann et al. (2024) find that by training an SAE on the decoder or another SAE, a technique called Meta-SAEs, it is possible to break down a single SAE latent like “Einstein” into subcomponents like “German” and “Physicist” and “starts with E”. Meta-SAEs may provide a promising future research direction to overcome the feature absorption phenomenon we describe in this paper.
@@ 6 Discussion
Interpretability of SAE latents
In this work, we use a simple first-letter identification task to investigate whether SAEs extract monosemantic and intrepretable features, and how this is affected by varying hyperparameters like SAE size, sparsity, or layer. We find that the SAE latents we investigated were not interpretable and that varying the sparsity or the size of the SAE did not meaningfully change this.
One may argue it is unreasonable to judge an SAE latent against a linear probe directly optimized to perform the same classification task and that it has been established that sometimes unsupervised methods can surprise us (Nanda, 2023). However, we argue that for a latent to be considered interpretable, its behavior should match what one would reasonably expect the latent to be doing after inspecting its activation patterns. In our experiments, we use SAE latents that do appear to perform first letter classification, and then evaluate how well they perform this task. We validate as well that these latents causally mediated model performance on the first-letter task in Appendix A.3. We are convinced that these latents should reasonably be considered “first-letter” latents and that their performance on the first-letter identification task is a valid measure of their interpretability.
Feature absorption
In trying to understand why SAE latents fail to match the performance of LR probes, we identified a form of feature splitting we call “feature absorption”. Feature absorption may be particularly problematic for SAEs because it creates an interpretability illusion where we believe we have found an interpretable latent, but absorption induces lower recall by creating clear false negatives to the mainline interpretation of the latent. For example, we may believe we have found a SAE latent which tracks deceptive behavior in the model, but due to feature absorption, there may be many cases where that latent fails to fire. This lower recall poses problems for methods which rely on using SAEs to find sparse circuits (Marks et al., 2024), as the number of latents needed to characterize model behavior may be much larger than expected. We find that feature absorption happens even in high-recall latents, so this is not only a problem for low L0 SAEs and appears to be a more fundamental issue.
Feature absorption is not predicted by toy-models where features do not co-occur (Bricken et al., 2023). While more work is required to understand why feature absorption occurs, we suspect it is a consequence of co-occurrence between sparse and dense features. If a dense feature like “starts with letter D” always co-occurs with a more sparse feature like “dogs”, the SAE can increase sparsity by absorbing the “starts with D” feature into a “dogs” latent.
It remains to be seen if we can predict or identify excepted instances where a feature “should have activated” but does not activate due to absorption. One promising direction is meta-SAEs, a novel method for decomposing SAE latents and may decompose absorbed features (Bussmann et al., 2024). One interpretation of our results is that competition may exist between “latents” and “meta-latents” for activation on particular examples and that re-allocation of examples between SAE latents enables SAEs to interpolate between different possible decompositions.
@@
A primary goal of future work should be to secure further external validity of our findings. This could include finding examples of feature absorption in SAEs trained on other models, with other architectures, or finding examples of feature absorption unrelated to character identification. We expect it should be possible to demonstrate feature absorption in a toy model by mixing dense features with sparse features that always co-occur with these dense features.
We hope as well this investigation may lead to research into solutions, particularly those involving Meta-SAEs (Bussmann et al., 2024), to solve or mitigate feature absorption. Anoter possible solution may be attribution dictionary learning (Olah et al., 2024a).
Limitations
The results presented in our paper are based on a single model (Gemma-2-2B) using one type of SAE architecture (JumpReLU). Our feature absorption metric requires having ground-truth knowledge of true labels to first train a LR probe, whereas many features of interest in a LLM lack such clear-cut ground-truth labels. Our metric uses ablation effect to ensure absorbed features causally mediate model behavior, but therefore cannot be easily used in final model layers.
7 Acknowledgements
@@ A.1 Glossary of Terms
When inserting a SAE into the computation path of the model, errors in SAE reconstruction will propagate to later parts of the model and can change the model output. We refer to the error as the SAE error term, and corresponds to the difference between the SAE output and the original SAE input activation. Marks et al. (2024) introduced the idea of adding this error term back to the SAE output to ensure that the SAE does not change model output.
@@
A platform, https://neuronpedia.org, which hosts feature dashboards for popular SAEs (Lin & Bloom, 2023).
@@
A phenomenon in SAEs introduced by Bricken et al. (2023), where a SAE latent tracks a general feature in a narrow SAE, but splits into multiple more specific SAE latents in a wider SAE. For instance, a latent tracking “starts with L” in a narrow SAE may split into a latent tracking “starts with capital L” and a latent tracking “starts with lowercase L” in a wider SAE.
@@
A problematic form of feature splitting where a SAE latent appears to track an interpretable feature, but that latent has seemingly arbitrary exception cases where it fails to fire. Instead, an approximately token-aligned feature “absorbs” the feature direction and fires in place of the main latent.
@@
In the context of neural network interpretability, a circuit refers to a subgraph of neurons or features within a neural network that work together to perform a specific function or computation. The study of circuits aims to understand how different components of a neural network interact to process information and produce outputs.
@@
A variant of linear probing where only the k most important features (as determined by some selection method) are used to train the probe. This helps identify which specific neurons or features are most relevant for a given task.
@@
An experimental method where a component of a system (in this case, a neuron or feature in a neural network) is removed or altered to observe its effect on the system’s performance. This helps determine the causal importance of the component.
@@
An attribution method that assigns importance scores to input features by accumulating gradients along a path from a baseline input to the actual input. In this paper, it’s used as an approximation technique for ablation studies.
@@
A.2 How good is Gemma-2 on character identification tasks?

### v5

@@ 6 Related work
Applications of Probes and SAEs for Model Interpretability
Probing methods can extract interpretable information from language models, though this does not guarantee the model uses these representations in its computation, and requires labeled data [7].
Prior work has shown that many human-interpretable concepts in LLM activations are represented as linear directions in activation space, known as the linear representation hypothesis [8, 28]. Li et al. [18] used non-linear probes to recover board representations from a transformer trained on Othello scripts (“OthelloGPT”). Nanda [23] later showed that linear representations were not only recoverable but also editable.
Karvonen et al. [16] developed objective metrics for SAE evaluation using Chess and Othello board states, but does not apply these to SAEs trained on LLM activations. Work by Olah et al. [26], Kissane et al. [17], Templeton et al. [31] noted poor precision/recall of SAE latents compared to known proxies. We extend this by showing how sparsity mediates precision/recall across many Gemma Scope SAEs and offer a possible explanation of low recall due to feature absorption.
Engels et al. [9] investigated SAE errors, finding that not all SAE error is linearly decomposable.
Studying precision and recall of SAE Latents
Most existing work on SAE interpretability mainly studies max activating examples [5], which may be misleading. There are more rigorous works which only measure precision [2, 31, 17]. Recent work has briefly explored recall and found it to be worse than expected naively, but this remains poorly understood [26]. We build on this work by evaluating precision / recall on a large number of SAEs, and offer a partial explanation for lower-than-expected recall of SAE latents in the form of feature absorption.
Decomposing SAE Latents
Feature splitting was first described in Bricken et al. [2], which noted that different SAE widths and sparsities induce latents of different granularity, with wider SAEs often learning more specific variants of features. Bussmann et al. [4] find that by training an SAE on the decoder of another SAE, a technique called Meta-SAEs, it is possible to break down a single SAE latent like “Einstein” into subcomponents like “German” and “Physicist” and “starts with E”.
@@ 7 Discussion
Limitations
Our Absoption metric uses ablation effect to ensure that the absorbed features causally mediate model behavior, and thus might not be easily transferable to the final model layers. Alternate metric formulations mitigating this are discussed in Appendix A.13. Due to compute constraints, we only train and evaluate a small number of non-JumpReLU SAEs in Figure 9. As our goal was only to show absorption occurs in all SAE architectures, we did not feel this is a significant drawback.
@@
The primary goal of future work is to find solutions to feature absorption. We are particularly hopeful that work extending Meta-SAEs [4] may solve or mitigate feature absorption. Another possible solution may be attribution dictionary learning [25]. Finally, structured sparsity techniques such as group lasso [13] or hierarchical sparse coding [14] may also be a promising direction of future work.
Other possible directions include allowing absorption to occur and using it as a way to recover hierarchies between features in a LLM. Our toy model results suggest that absorption leads to an asymmetric pattern in the encoder and decoder of the SAE, so it may be possible to use this insight to detect absorption (although there may be other reasons for an asymmetry in the SAE encoder and decoder beyond absorption).
Conclusion
We identify a form of feature splitting we call “feature absorption”, where more specific latents “steal credit” from more general ones. Absorption creates an interpretability illusion, where a seemingly interpretable latent has arbitrary false negatives in its mainline interpretation. Lower recall poses problems for using SAEs for high-stakes classification or finding sparse circuits [21], as the number of latents needed to characterize model behavior may be much larger than expected.
We show that absorption is a consequence of hierarchical co-occurrence between sparse and dense features. If a dense feature like “starts with letter D” always co-occurs with a more sparse feature like “dogs”, the SAE can increase sparsity by absorbing the “starts with D” feature into a “dogs” latent.
We hope that our work highlights the fundamental limitations of sparse feature extraction and prompts future research on SAEs such as identifying cases where a feature “should have activated” but does not due to absorption, and exploring theoretical solutions to absorption. The ease of demonstrating absorption in toy models makes it easier to validate potential solutions.
Acknowledgments and Disclosure of Funding
@@ A.1 Glossary of Terms
When inserting a SAE into the computation path of the model, errors in SAE reconstruction will propagate to later parts of the model and can change the model output. We refer to the error as the SAE error term, and corresponds to the difference between the SAE output and the original SAE input activation. Marks et al. [21] introduced the idea of adding this error term back to the SAE output to ensure that the SAE does not change model output.
@@
A platform, https://neuronpedia.org, which hosts feature dashboards for popular SAEs [20].
Token-aligned latent:
A latent which seems to roughly fire on variants of the same token. For instance, a “Snake” token-aligned latent may fire on the tokens “Snake”, “SNAKE”, “_snakes”, etc…
@@
A phenomenon in SAEs introduced by Bricken et al. [2], where a SAE latent tracks a general feature in a narrow SAE, but splits into multiple more specific SAE latents in a wider SAE. For instance, a latent tracking “starts with L” in a narrow SAE may split into a latent tracking “starts with capital L” and a latent tracking “starts with lowercase L” in a wider SAE.
@@
A problematic form of feature splitting where a SAE latent appears to track an interpretable feature, but that latent has seemingly arbitrary exception cases where it fails to fire. Instead, a more specific latent “absorbs” the feature direction and fires in place of the main latent.
@@
In the context of neural network interpretability, a circuit refers to a subgraph of neurons or latents within a neural network that work together to perform a specific function or computation. The study of circuits aims to understand how different components of a neural network interact to process information and produce outputs.
@@
A variant of linear probing where only the k most important latents (as determined by some selection method) are used to train the probe. This helps identify which specific neurons or latents are most relevant for a given task.
@@
An experimental method where a component of a system (in this case, a neuron or latent in a neural network) is removed or altered to observe its effect on the system’s performance. This helps determine the causal importance of the component.
@@
An attribution method that assigns importance scores to input latents by accumulating gradients along a path from a baseline input to the actual input. In this paper, it’s used as an approximation technique for ablation studies.
@@
A.2 Proof: absorption decreases SAE loss for hierarchical features
We analyze the effect of a specific form of feature absorption, termed $\delta$-absorption, within a Sparse Autoencoder (SAE) framework. We consider two hierarchically related features, $f_{1}$ and $f_{2}$ (where $f_{2}\subset f_{1}$), and demonstrate that for a defined family of encoder and decoder weights parameterized by $\delta\in[0,1]$ ($\delta=0$ corresponds to no absorption, and $\delta=1$ corresponds to full absorption):
  1. 1.
    Perfect reconstruction of inputs composed of these features is maintained across all values of $\delta$.
  2. 2.
    The sparsity loss component attributable to these features is a decreasing function of $\delta$, provided the child feature $f_{2}$ has a non-zero probability of appearing.
  3. 3.
    Consequently, optimizing for sparsity encourages higher values of $\delta$, i.e., greater absorption.
1. Preliminaries and Assumptions
H1. Dataset and Features
Let $\mathcal{D}$ be a dataset. We consider a set of features =$\mathcal{F}=\{f_{1},f_{2},\ldots,f_{d}\}$. Each feature $f_{i}\in\mathbb{R}^{k}$ is a vector with unit norm, $\|f_{i}\|_{2}=1$. Features are mutually orthogonal: $f_{i}\cdot f_{j}=\delta_{ij}$ (Kronecker delta), where $\delta_{ij}=1$ if $i=j$ and 00 if $i\neq j$. An activation $h\in\mathbb{R}^{k}$ in the model’s residual stream is a linear combination of active features: $h=\sum_{j\in\text{ActiveFeatures}}f_{j}$.
H2. Feature Hierarchy and Probabilities
We focus on two features $f_{1},f_{2}\in\mathcal{F}$ with a hierarchy $f_{2}\subset f_{1}$. This implies that if $f_{2}$ is present in a datapoint, $f_{1}$ must also be present. The probabilities of observing combinations of $f_{1}$ and $f_{2}$ are:
  • $p(f_{1},f_{2})=p_{11}$: Probability of $f_{1}$ and $f_{2}$ co-occurring (e.g., input is $f_{1}+f_{2}$).
  • $p(f_{1},\neg f_{2})=p_{10}$: Probability of $f_{1}$ occurring without $f_{2}$ (e.g., input is $f_{1}$).
  • $p(\neg f_{1},f_{2})=p_{01}$: Probability of $f_{2}$ occurring without $f_{1}$. By the hierarchy assumption, $p_{01}=0$.
  • $p(\neg f_{1},\neg f_{2})=p_{00}$: Probability of neither $f_{1}$ nor $f_{2}$ occurring (e.g., input is 00 or some $f_{k},k\neq 1,2$).
We assume $p_{11}+p_{10}+p_{00}=1$.
H3. Sparse Autoencoder (SAE) Model
The SAE reconstructs an input $h$ as $\hat{h}=f_{\phi}(h)$. The reconstruction is $\hat{h}=W_{d}z$, where $z=\text{ReLU}(W_{e}h)$. No bias terms are used. $W_{e,i}$ is the $i$-th row of $W_{e}$ (encoder vector for latent $i$), and $W_{d,i}$ is the $i$-th column of $W_{d}$ (decoder vector for latent $i$). We analyze two specific latents, $z_{1}$ and $z_{2}$, intended to capture $f_{1}$ and $f_{2}$. Other latents $z_{j}$ for $j>2$ are assumed to perfectly reconstruct other features $f_{j}$ (e.g. $W_{e,j}=f_{j},W_{d,j}=f_{j}$) and do not interact with $f_{1},f_{2}$ due to orthogonality.
H4. SAE Loss Function
The total loss is $\mathcal{L}=\mathcal{L}_{\text{rec}}+\lambda\mathcal{L}_{\text{sp}}$, where $\lambda>0$. $\mathcal{L}_{\text{rec}}=\mathbb{E}_{h\sim\mathcal{D}}\left[\|h-\hat{h}\|_{2}^{2}\right]$. $\mathcal{L}_{\text{sp}}=\mathbb{E}_{h\sim\mathcal{D}}\sum_{i}|z_{i}|$. Our analysis will focus on the contributions of $z_{1},z_{2}$ to $\mathcal{L}_{\text{rec}}$ and $\mathcal{L}_{\text{sp}}$.
H5. Definition of $\delta$-Absorption
We define a specific parameterization for the encoder and decoder weights associated with $f_{1}$ and $f_{2}$ by a parameter $\delta\in[0,1]$:
  • $W_{e,1}=f_{1}-\delta f_{2}$
  • $W_{e,2}=f_{2}$
  • $W_{d,1}=f_{1}$
  • $W_{d,2}=f_{2}+\delta f_{1}$
$\delta=0$ represents no absorption, while $\delta=1$ represents full absorption.
2. Proposition 1: Perfect Reconstruction under $\delta$-Absorption
For any $\delta\in[0,1]$, and for inputs $h$ consisting only of $f_{1},f_{2}$ or 00, the reconstruction $\hat{h}=W_{d,1}z_{1}+W_{d,2}z_{2}$ perfectly reconstructs $h$, i.e., the reconstruction loss component $\mathcal{L}_{\text{rec}}^{(1,2)}$ due to these features is 00.
Proof
We consider the possible input types based on $f_{1},f_{2}$:
Case 1: $h=f_{1}$ (only parent feature $f_{1}$ is present).
The latent activations are:
$z_{1}$ $=\text{ReLU}(W_{e,1}\cdot h)=\text{ReLU}((f_{1}-\delta f_{2})\cdot f_{1})$
$=\text{ReLU}(f_{1}\cdot f_{1}-\delta f_{2}\cdot f_{1})$
[…]
The reconstruction is:
$\hat{h}=z_{1}W_{d,1}+z_{2}W_{d,2}=1\cdot f_{1}+0\cdot(f_{2}+\delta f_{1})=f_{1}$
Thus, $\hat{h}=h$.
Case 2: $h=f_{1}+f_{2}$ (both parent $f_{1}$ and child $f_{2}$ are present).
The latent activations are:
$z_{1}$ $=\text{ReLU}(W_{e,1}\cdot h)=\text{ReLU}((f_{1}-\delta f_{2})\cdot(f_{1}+f_{2}))$
$=\text{ReLU}(f_{1}\cdot f_{1}+f_{1}\cdot f_{2}-\delta f_{2}\cdot f_{1}-\delta f_{2}\cdot f_{2})$
[…]
Since $\delta\in[0,1]$, $1-\delta\geq 0$, so $z_{1}=1-\delta$.
$z_{2}$ $=\text{ReLU}(W_{e,2}\cdot h)=\text{ReLU}(f_{2}\cdot(f_{1}+f_{2}))$
$=\text{ReLU}(f_{2}\cdot f_{1}+f_{2}\cdot f_{2})$
[…]
The reconstruction is:
$\hat{h}$ $=z_{1}W_{d,1}+z_{2}W_{d,2}=(1-\delta)f_{1}+1\cdot(f_{2}+\delta f_{1})$
$=(1-\delta)f_{1}+f_{2}+\delta f_{1}=f_{1}-\delta f_{1}+f_{2}+\delta f_{1}=f_{1}+f_{2}$
Thus, $\hat{h}=h$.
Case 3: $h=0$ (neither $f_{1}$ nor $f_{2}$ is present).
$z_{1}$ $=\text{ReLU}((f_{1}-\delta f_{2})\cdot 0)=0$
$z_{2}$ $=\text{ReLU}(f_{2}\cdot 0)=0$
The reconstruction is:
$\hat{h}=0\cdot f_{1}+0\cdot(f_{2}+\delta f_{1})=0$
Thus, $\hat{h}=h$.
Case 4: $h=f_{2}$ (only child feature $f_{2}$ is present).
This case is disallowed by assumption H2 ($p_{01}=0$), as $f_{2}\subset f_{1}$ implies $f_{1}$ must be present if $f_{2}$ is.
In all permissible cases, $h-\hat{h}=0$, so $\|h-\hat{h}\|_{2}^{2}=0$. Therefore, the reconstruction loss component due to $f_{1},f_{2}$, denoted $\mathcal{L}_{\text{rec}}^{(1,2)}$, is 0 for any $\delta\in[0,1]$.
3. Proposition 2: Sparsity Loss under $\delta$-Absorption
The expected sparsity loss contribution from latents $z_{1}$ and $z_{2}$, denoted $\mathcal{L}_{\text{sp}}^{(1,2)}=\mathbb{E}_{h\sim\mathcal{D}}[|z_{1}|+|z_{2}|]$, is given by:
$\mathcal{L}_{\text{sp}}^{(1,2)}=p_{11}(2-\delta)+p_{10}$
Furthermore, its derivative with respect to $\delta$ is:
$\frac{d\mathcal{L}_{\text{sp}}^{(1,2)}}{d\delta}=-p_{11}$
Proof
We calculate the sum of absolute latent activations $|z_{1}|+|z_{2}|$ for each case from Proposition 1 and weight them by their probabilities (H2):
  • If $h=f_{1}+f_{2}$ (probability $p_{11}$): $z_{1}=1-\delta$, $z_{2}=1$. Since $\delta\in[0,1]$, $1-\delta\geq 0$, so $|z_{1}|=1-\delta$. $|z_{2}|=1$. Thus, $|z_{1}|+|z_{2}|=(1-\delta)+1=2-\delta$.
  • If $h=f_{1}$ (probability $p_{10}$): $z_{1}=1$, $z_{2}=0$. Thus, $|z_{1}|+|z_{2}|=1+0=1$.
  • If $h=0$ (probability $p_{00}$, neither $f_{1}$ nor $f_{2}$ present): $z_{1}=0$, $z_{2}=0$. Thus, $|z_{1}|+|z_{2}|=0+0=0$.
The case corresponding to $p_{01}$ does not occur.
The expected sparsity loss from $z_{1},z_{2}$ is:
$\mathcal{L}_{\text{sp}}^{(1,2)}$ $=p_{11}\cdot(2-\delta)+p_{10}\cdot 1+p_{00}\cdot 0$
$=p_{11}(2-\delta)+p_{10}$
Taking the derivative with respect to $\delta$:
$\frac{d\mathcal{L}_{\text{sp}}^{(1,2)}}{d\delta}=\frac{d}{d\delta}(2p_{11}-\delta p_{11}+p_{10})$
Since $p_{11}$ and $p_{10}$ are constants with respect to $\delta$:
$\frac{d\mathcal{L}_{\text{sp}}^{(1,2)}}{d\delta}=-p_{11}$
$\square$
4. Corollary: Increasing Absorption Decreases Sparsity Loss
If $p_{11}>0$ (i.e., the child feature $f_{2}$ co-occurs with $f_{1}$ with non-zero probability), then increasing $\delta$ strictly decreases $\mathcal{L}_{\text{sp}}^{(1,2)}$.
Proof
From Proposition 2, $\frac{d\mathcal{L}_{\text{sp}}^{(1,2)}}{d\delta}=-p_{11}$. If $p_{11}>0$, then $-p_{11}<0$. A negative derivative implies that $\mathcal{L}_{\text{sp}}^{(1,2)}$ is a decreasing function of $\delta$ for $\delta\in[0,1]$. The minimum value of $\mathcal{L}_{\text{sp}}^{(1,2)}$ over this interval occurs at $\delta=1$ (full absorption), yielding $\mathcal{L}_{\text{sp}}^{(1,2)}(\delta=1)=p_{11}(2-1)+p_{10}=p_{11}+p_{10}$. The maximum value occurs at $\delta=0$ (no absorption), yielding $\mathcal{L}_{\text{sp}}^{(1,2)}(\delta=0)=p_{11}(2-0)+p_{10}=2p_{11}+p_{10}$. $\square$
5. Conclusion
Given the specified $\delta$-absorption mechanism for an SAE handling two hierarchical features $f_{1},f_{2}$ (where $f_{2}\subset f_{1}$):
  1. 1.
    Perfect reconstruction of inputs composed of $f_{1}$ and $f_{2}$ is maintained irrespective of the degree of absorption $\delta$. Thus, $\mathcal{L}_{\text{rec}}^{(1,2)}$ is unaffected by $\delta$.
  2. 2.
    The sparsity loss component $\mathcal{L}_{\text{sp}}^{(1,2)}$ associated with these features is $p_{11}(2-\delta)+p_{10}$.
  3. 3.
    If $p_{11}>0$, the total loss $\mathcal{L}$ (focusing on the components related to $f_{1},f_{2}$) decreases as $\delta$ increases because $\mathcal{L}_{\text{rec}}^{(1,2)}$ is constant (zero) and $\mathcal{L}_{\text{sp}}^{(1,2)}$ decreases.
  4. 4.
    Therefore, an optimization process like gradient descent, when minimizing the total loss $\mathcal{L}=\mathcal{L}_{\text{rec}}+\lambda\mathcal{L}_{\text{sp}}$ (where $\lambda>0$), will favor increasing $\delta$ towards 1, thereby promoting feature absorption for these hierarchically related features, assuming the SAE learns or is constrained to these forms of $W_{e}$ and $W_{d}$.
This formalizes the argument that, under the given conditions and definitions, absorption is a mechanism that can reduce SAE loss by improving sparsity without harming reconstruction for hierarchical features.
A.3 Extended toy model experiments
In this section we explore further variants on absorption in toy models. We use the same setting as our main toy model experiment, with four mutually-orthogonal true features, and train an SAE with four latents. Each true feature $f\in\mathbb{R}^{50}$. Unless otherwise stated, every time feature 1 fires feature 0 must also fire, but feature 0 is allowed to fire on its own as well. This is to simulate hierarchal features such as our example “starts with S” and “short” features, where every time the “short” feature fires we expect “starts with S” must also fire since “short” starts with “S”, but “starts with S” can fire on its own as well. Feature 2 and 3 are fully independent. All features fire with magnitude 1.0 and variance 0.0 unless otherwise stated.
Magnitude variance causes partial absorption
In our main toy model experiment, each true feature fires with magnitude exactly 1.0. This is not very realistic, though - likely there will be some variance in feature firing magnitudes in real LLMs. We simulate this by adding variance of 0.1 to the firing magnitude of feature 0, so the relative magnitudes of feature 0 and 1 are no longer fixed. We show the plots of cosine similarity between SAE encoder and decoder in Figure 10. Here, we still see the same absorption pattern in the SAE encoder and decoder with the latent 3 encoder containing a negative component of feature 1, and the latent 0 decoder merging features 0 and 1. We show some sample true feature firings and corresponding SAE latent activations in Table 2.
Figure 10: SAE encoder and decoder with true features. The firing magnitude of feature 0 has variance 0.1, while the remaining features fire with variance 0.0. When there is variance in the firing magnitudes of parent and child features, we still see an absorption pattern in the SAE encoder and decoder with the latent 3 encoder containing a negative component of feature 1, and the latent 0 decoder merging features 0 and 1.
Table 2: Sample true feature firings and corresponding SAE latent activations. Feature 1 only fires if feature 0 fires. Feature 0 has variance of 0.1 in its firing magnitude, while the other feature have no variance in their firing magnitude.
True features SAE Latent acts
1.00 0.00 0.00 0.00 0.00 0.00 0.00 1.00
1.00 1.00 0.00 0.00 1.27 0.00 0.00 0.22
0.90 1.00 0.00 0.00 1.27 0.00 0.00 0.12
0.75 1.00 0.00 0.00 1.26 0.00 0.00 0.00
We see that now the SAE latent tracking feature 0 still fires when the true values of features 0 and 1 are both 1.0, but very weakly. However, if the magnitude of feature 0 drops down to 0.75, then the feature 0 latent fully turns off.
We call this phenonemon partial absorption. In partial absorption, there’s co-occurrence between a dense and sparse feature, and the sparse feature absorbs the direction of the dense feature. However, the SAE latent tracking the dense feature still fires when both the dense and sparse feature are active, only very weakly. If the magnitude of the dense feature drops below some threshold, it stops firing entirely.
Feature absorption is an optimal strategy for minimizing the L1 loss and maximizing sparsity. However, when a SAE absorbs one latent into another, the absorbing latent loses the ability to modulate the magnitudes of the underlying features relative to each other. The SAE can address this by firing the latent tracking the dense feature as a "correction" to add back some of the dense feature direction into the reconstruction. Since the dense feature latent is firing weakly, it still has lower L1 loss than if the SAE fully separated out the features into their own latents.
Imperfect co-occurrence can still lead to partial absorption
Next, we test what will happen if feature 1 is more likely to fire if feature 0 is active, but can still fire without feature 0. We set up feature 1 to fire with feature 0 95% of the time, but 5% of the time it can fire on its own. For this experiment, all features fire with magnitude 1.0 and 0 variance. We show the cosine similarities of the SAE encoder and decoder with true features in Figure 11. Some sample feature firings and corresponding SAE activations are shown in Table 3.
Figure 11: SAE encoder and decoder with true features. Feature 1 fires with feature 0 95% of the time, but 5% of the time feature 1 is allowed to fire on its own. Even though there is not perfect co-occurrence between features 0 and 1, we still see the SAE encoder and decoder learn a weak absorption pattern, with the decoder latent for feature 1 absorbing part of feature 0, and the encoder latent for feature 0 including a negative component of feature 1.
Table 3: Sample feature values and corresponding SAE activations. Feature 1 can only fire if feature 0 is active 95% of the time, but 5% of the time feature 1 can fire on its own. We see signs of partial absorption, where the latent tracking feature 0 fires noticeably more weakly if feature 1 is active.
True features SAE Latent acts
1.00 0.00 0.00 0.00 0.00 0.00 0.00 1.00
1.00 1.00 0.00 0.00 0.00 1.05 0.00 0.67
0.00 1.00 0.00 0.00 0.00 0.95 0.00 0.00
We see signs of partial absorption here as well. We see the same absorption pattern in the SAE encoder and decoder as we saw in our other absorption examples, although less severe than the previous examples. We also see in the sample firing patterns that when both feature 0 and 1 fire together, the latent tracking feature 0 fires with noticeably lower magnitude than when feature 0 fires on its own. Here, even though the co-occurrence between features 0 and 1 is not perfect, we still see partial absorption.
Absorption also affects TopK SAEs
So far, we have only shown feature absorption occurring with standard L1 SAEs. Next, we examine how other absorption affects other architectures using a batch topk SAE [3]. Batch topk SAEs are an improved version of topk SAEs [10] where the top $k*B$ latents are used to reconstruct the SAE input, where $B$ is the batch size. As the topk function enforces sparsity, there is no additional L1 loss term.
Topk SAEs are harder to use for very small toy models like our 4-feature toy model above, since if the $k$ is too large relative to the size of the SAE the SAE will not learn correct features. To address this, we use a slightly larger toy model with 12 mutually orthogonal true features. All features fire independently with probability 0.15, except for the first 2 features. Feature 0 is the parent feature in our hierarchy, and fires with probability 0.4. Feature 1 is the child feature, and fires with probability 0.6 only if feature 1 fires, but never fires if feature 1 does not fire. All features fire with magnitude 1.0. We train a batch topk SAE with $k=2$. We show the cosine similarities of the SAE encoder and decoder with true features in Figure 12.
Figure 12: SAE encoder and decoder with true features for a batch topk SAE with k=12. Feature 1 is only allowed to fire if feature 0 fires. We still see a very clear absorption pattern between the latents tracking features 0 and 1 despite the lack of L1 loss.
We still see a clear absorption pattern between the latents tracking features 0 and 1 despite the lack of L1 loss. Absorption increases sparsity, which allows the topk SAE to have better reconstruction loss at a given k, and is thus what the SAE learns.
A.4 Ablation algorithm
Algorithm 1 SAE Latent Ablation
  Insert SAE in model computation, including error term
  Define a scalar metric on the model’s output distribution
  Calculate baseline metric value for a test prompt
  for each token of interest do
     for each SAE latent do
        Set the SAE latent activation to 0
        Recalculate the metric
        Compute ablation effect (baseline - new metric)
        Reset the SAE latents to its original value
     end for
  end for
A.5 How good is Gemma-2 on character identification tasks?